OpenStudy (anonymous):

An angle in the standard position of 201.8 degrees has a point P= (x,-2) on the terminal side. Find the value of x.

OpenStudy (anonymous):

@fprt Simply put \[\huge \tan\left(201.8^o\right)=\frac{y}{x}\]

OpenStudy (anonymous):

So, transposing, so you the value of x \[\huge x=\frac{y}{\tan201.8^o}\]

OpenStudy (anonymous):

hey thanks soo much! i think the answer i got is correct.

OpenStudy (anonymous):

That's good :)

OpenStudy (anonymous):

I have another question: Find a positive value in the equation 1+3cos pi/8(x+7)=0 How would I go about approaching this problem?

OpenStudy (anonymous):

@PeterPan

OpenStudy (anonymous):

I don't get your question :(

OpenStudy (anonymous):

Well i dont understand the question either, it is question 4 in the attached file

OpenStudy (anonymous):

I really do not understand :(

OpenStudy (anonymous):

oh ok thanks for your help